he percentage of microorganisms remaining vs. exposure time to autoclave sterilization process follows % = 70-t*2. The percentage of microorganisms remaining vs. exposure time to dry heat sterilization process follows % = 80-t*5. At what time both the methods are equally efficient?
After 6 mins of sterilization what will be the difference between the percentages of microorganisms remaining from both these methods, give the absolute number
1
Expert's answer
2020-04-24T13:38:50-0400

As per the question requirement for both to be equally efficient

70-2t=80-5t

3t=10 t=3.33

So, after t=3.33 both will be equally efficient.

Now we would consider individually

% autoclave after 6 min= 70-6"\times2=70-12=58 %" %

and for dry heat asfter 6 min = "80 -5\times6=50" %

so here difference will be = 58-50 = 8 %